# CodeBrain: Bridging Decoupled Tokenizer and Multi-Scale Architecture for EEG Foundation Model

### v1

@@ 3.3 TFDual-Tokenizer Pretraining
where $\mathrm{sim}(\mathbf{h}_{i},\mathbf{h}_{j})$ is the cosine similarity, $\tau>0$ is the temperature parameter, and $\mathbb{1}_{[k\neq i]}$ is an indicator function used to exclude itself. We introduce a T-Decoder to reconstruct raw signals from the temporal code embedding $\tilde{e}_{n}$, which consists of a Transformer encoder $TF^{(t)}$ followed by a linear projection. Let $y_{t}$ be the Transformer output from Eq. (2); the overall training objective is:
$\mathcal{L}_{t}=\mathcal{L}_{CL}+||y_{t}-x_{t}||^{2}_{2},\quad y_{t}=TF^{(t)}(MLP(\tilde{e}_{n})).$ (9)
Finally, the training objective for the TFDual-Tokenizer is defined as:
@@
$\mathcal{L}_{codebook}=\sum_{x\in\mathcal{D}}\sum_{i}\underbrace{||y_{n}^{A}-A_{n}||^{2}_{2}+||y_{n}^{P}-\phi_{n}||^{2}_{2}}_{\text{reconstruction loss}}+\underbrace{\|sg(\tilde{e}_{n})-v_{j}^{(t)}\|^{2}+\|sg(\tilde{e}_{n})-v_{j}^{(f)}\|^{2}}_{\text{codebook loss}}+\mathcal{L}_{t},$ (10)
@@
where $sg(\cdot)$ denotes stop-gradient, avoiding the update of encoder parameters.
@@ 3.4 EEGSSM Pretraining
In this stage, we introduce a novel convolutional structured state space model framework, called EEGSSM, composed of multiple EEGSSM blocks. First, raw EEG signals are projected into a higher-dimensional embedding space via a 1D convolutional layer. The resulting features are processed by EEGSSM blocks. Another 1D convolutional layer then maps the output back to the token space for reconstructing the indices of masked tokens produced by the TFDual-Tokenizer.
EEGSSM Block Our EEGSSM block is composed of several blocks, which are integrated together through a residual connection mechanism. An EEGSSM block consists of an RMS Norm layer, SGConv layer, SWA layer, and a gating component. Root Mean Square Layer Normalization (RMSNorm) zhang2019rootmeansquarelayer can be considered an improved version of LayerNorm, normalizing by using the square root of the mean to reduce the impact of noise. Afterward, we feed the intermediate variables into the SGConv layer to obtain a global receptive field through convolution SSM.
SGConv Layer. SGConv is a structured SSM model using convolution architecture (See supplemental material), and its convolution structure can be represented as a FFT formula:
@@
where $F_{N}$ denotes the DFT matrix of size $N$. This FFT convolution has a computational complexity of $O(nlog(n))$. As a type of convolutional SSM, SGConv improves the convolution kernel $\overline{K}$ in Eq. (11) by introducing two features: sparse parameterization and kernel decay, making SGConv easier and more efficient to compute compared to the traditional S4 kernel. Let $L$ be the length of the input sequence. The convolution kernel $\overline{K}$ of SGConv is composed of several sub-kernels. Assuming the size of the first sub-kernel is $d$, with parameters $w_{i}\in\mathbb{R}^{d\times d}$, then the number of sub-kernels can be expressed as $N=\log_{2}(\frac{L}{d})+1$. The convolution kernel $\overline{K}$ in Eq. (11) can thus be initialized as:
@@
where $\alpha$ denotes the decay coefficient, usually chosen to be 1/2, inducing the decaying structure, and $\text{Upsample}(x)$ denotes upsample $x$ to length $l$. We also introduce the normalization constant $Z$ to ensure that the convolution operation does not change the scale of the input.
@@
At the end of the EEGSSM block, we use a gating mechanism to control the output of the block. We employ a gated unit similar to Wavenet obeid2016temple , which can suppress useless or irrelevant features and help stabilize training in deep networks. We concatenate the global features output $y_{sg}$ by the SGConv layer with the local features output $y_{swa}$ by SWA and feed them into a gated unit:
@@
where $tanh(\cdot)$ and $\sigma(\cdot)$ are the tanh function and sigmoid function, $\odot$ denotes an element-wise multiplication operator, $W_{f}$ and $W_{g}$ are learnable convolution filters. $y_{1}$ becomes the input to the next block, while $y_{2}$ will be aggregated to the output of SSM blocks through a skip connection.
Pre-Training Objective To help the EEGSSM model learn general EEG representations, we use a Masked Autoencoder (MAE) for self-supervised pre-training. For a patched sample $X=\{x_{i}\mid i\in[1,2,\ldots,C]\}$, we randomly generate a mask =$\mathcal{M}=\{m_{i}\mid i\in[1,2,\ldots,C]\}$ from a Bernoulli distribution of $r$ proportion, where $m_{ij}\in\{0,1\}$. We reconstruct the token indices of masked EEG patches from the TFDual-Tokenizer by cross-entropy loss. Let $y_{i}$ denote the output of the EEGSSM block, the probability that the EEG signal matches the corresponding token $v_{i}$ in the codebooks:
@@ 4.1 Datasets
Pre-Training We pretrain CodeBrain on the TUH EEG Corpus (v2.0.1, modified in March 2025) obeid2016temple , the largest publicly available EEG dataset to date. Data processing follows a standardized pipeline: recordings shorter than 5 minutes are excluded, and the first and last minute of each segment is removed. We retain 19 commonly used EEG channels (C3, C4, Cz, F3, F4, Fp1, Fp2, F7, F8, Fz, O1, O2, P3, P4, Pz, T3, T4, T5, T6), selected based on the international 10–20 system for electrode placement hh1958ten , apply band-pass filtering (0.3–75 Hz), and notch filtering at 60 Hz to remove noise. The data is resampled to 200 Hz and divided into 30-second non-overlapping segments. Segments with absolute amplitudes over $100\leavevmode\nobreak\ \mu\mathrm{V}$ are filtered out. To normalize the signals, each value is divided by 100. Each segment is split into 1-second windows, resulting in 570 EEG patches per sample. After preprocessing, 1,109,545 samples (about 9,246 hours) are retained for pretraining.
Downstream Tasks. We evaluate CodeBrain on 8 downstream tasks across 10 publicly available EEG datasets, covering a diverse range of applications. Detailed dataset configurations are in Table 1. We perform cross-subject or cross-session splits with strict separation between training, validation, and test sets. For FACED chen2023large , we use 80 subjects for training, 20 for validation, and the remaining 23 for testing. In SEED-V liu2021comparing , we perform a leave-one-session-out manner. ISRUC_S3 khalighi2016isruc consists of 10 subjects, for which we apply an 8:1:1 cross-subject split. MentalArithmetic consists of 36 subjects, and we use a 7:1:1 cross-subject split. BCIC2020-T3 jeong20222020 follows the official competition protocol. In CHB-MIT shoeb2009application , we use recordings from 19 subjects for training, and from 2 subjects each for validation and testing. Additional dataset details are provided in the supplementary material.
@@
Downstream Tasks Datasets Shape #Samples Class
Emotion Recognition FACED chen2023large 32 channels * 10s 10,332 9-class
SEED-V liu2021comparing 62 channels * 1s 117,744 5-class
Sleep Staging ISRUC_S1 khalighi2016isruc 6 channels * 30s 86,320 5-class
ISRUC_S3 khalighi2016isruc 6 channels * 30s 8,500 5-class
Imagined Speech Classification BCIC 2020-T3 jeong20222020 64 channels * 3s 6,000 5-class
Mental Stress Detection Mental Arithmetic mumtaz2016eeg 20 channels * 5s 1,707 2-class
Seizure Detection CHB-MIT shoeb2009application 16 channels * 10s 326,993 2-class
Motor Imagery Classification SHU-MI goldberger2000physiobank 32 channels * 4s 11,988 2-class
Event Type Classification TUEV obeid2016temple 16 channels * 5s 112,491 6-class
Abnormal Detection TUAB obeid2016temple 16 channels * 10s 409,455 2-class
4.2 Experiment Settings
Two-Stage Training Setup. All experiments are conducted on NVIDIA 40GB A100 GPUs. In Stage 1, the TFDual-Tokenizer is trained with temporal and frequency codebooks of 4096 codes (32 dimensions) for 20 epochs, using a batch size of 256 and a learning rate of 1e-4, across six A100 GPUs for approximately ten hours. In Stage 2, an 8-layer EEGSSM backbone with a masking ratio of 0.5 is trained for 10 epochs, using a batch size of 256 on two A100 GPUs for about one day.
Fintuning Strategy. We adopt a linear probing setup to evaluate the quality of pretrained representations. All downstream task datasets are resampled to 200 Hz to match the pretraining configuration. A three-layer MLP is applied. The first layer aggregates information across all EEG channels. The second layer compresses the entire $x$-second sequence (depending on task configuration) into a 200-dimensional vector. The final layer maps this representation to the target class. Each hidden layer uses ELU activation and dropout to improve generalization.
Baselines. We compare our model with five publicly available EFMs that have released pretrained weights, representing different pretraining strategies. BENDR kostas2021bendr adopts contrastive learning to learn general EEG representations. BIOT yang2023biot uses patch-based continuous tokenization. LaBraM introduces discrete neural tokens through vector quantization jianglarge . EEGPT wang2024eegpt and CBraMod wang2024cbramod both rely on masked reconstruction of raw EEG signals.
4.3 Comparison with Baselines
We ensure consistent data splits across all baselines. Results are reported on six representative downstream datasets, with additional results provided in the supplementary material. As shown in Table 2, CodeBrain achieves consistent performance gains compared to baselines. For multi-class classification, it achieves the largest gain of +0.0911 in Cohen’s Kappa (21.6%), +0.0718 in Weighted F1 score (13.3%), +0.0728 in Balanced Acc (13.5%) on BCIC 2020-T3 over the strongest baseline wang2024cbramod . For binary classification, it achieves the largest gain of +0.0802 in AUROC (10.1%), +0.0910 in AUC-PR (14.5%) and +0.0258 in Balanced Acc (3.6%) on Mental Arithmetic over the strongest baseline wang2024cbramod . These results demonstrate the superior generalizability of CodeBrain.
@@ 4.3 Comparison with Baselines
Kappa W-F1 Bal. Acc Kappa W-F1 Bal. Acc
BENDRkostas2021bendr 0.4716 ± 0.0095 0.5340 ± 0.0086 0.5320 ± 0.0083 0.0335 ± 0.0062 0.2026 ± 0.0330 0.2231 ± 0.0059
BIOTyang2023biot 0.4476 ± 0.0254 0.5136 ± 0.0112 0.5118 ± 0.0118 0.2261 ± 0.0262 0.3856 ± 0.0203 0.3837 ± 0.0187
LaBraMjianglarge 0.4698 ± 0.0102 0.5288 ± 0.0188 0.5273 ± 0.0107 0.2386 ± 0.0209 0.3974 ± 0.0111 0.3976 ± 0.0138
EEGPTwang2024eegpt 0.4639 ± 0.0023 0.3924 ± 0.0017 0.4607 ± 0.0014 0.1323 ± 0.0062 0.3090 ± 0.0052 0.3061 ± 0.0044
CBraModwang2024cbramod 0.5041 ± 0.0122 0.5618 ± 0.0093 0.5509 ± 0.0089 0.2569 ± 0.0143 0.4101 ± 0.0108 0.4091 ± 0.0097
CodeBrain 0.5270 ± 0.0039 0.5849 ± 0.0033 0.5816 ± 0.0028 0.2735 ± 0.0032 0.4235 ± 0.0022 0.4137 ± 0.0023
@@
Kappa W-F1 Bal. Acc Kappa W-F1 Bal. Acc
BENDRkostas2021bendr 0.5995 ± 0.0151 0.6789 ± 0.0142 0.6352 ± 0.0095 0.0607 ± 0.0093 0.2379 ± 0.0165 0.2485 ± 0.0075
BIOTyang2023biot 0.7168 ± 0.0119 0.7834 ± 0.0096 0.7598 ± 0.0109 0.3650 ± 0.0176 0.4917 ± 0.0079 0.4920 ± 0.0086
LaBraMjianglarge 0.7194 ± 0.0162 0.7843 ± 0.0189 0.7617 ± 0.0122 0.3800 ± 0.0242 0.5054 ± 0.0205 0.5060 ± 0.0155
EEGPTwang2024eegpt 0.6160± 0.0856 0.6375 ± 0.0632 0.6650 ± 0.0311 0.0567 ± 0.0164 0.2441 ± 0.0105 0.2453 ± 0.0131
CBraModwang2024cbramod 0.7407 ± 0.0251 0.8056 ± 0.0219 0.7844 ± 0.0126 0.4216 ± 0.0163 0.5383 ± 0.0096 0.5373 ± 0.0108
@@
AUROC AUC-PR Bal. Acc AUROC AUC-PR Bal. Acc
BENDRkostas2021bendr 0.6248 ± 0.0765 0.3661 ± 0.0672 0.5681 ± 0.0448 0.8632 ± 0.0526 0.3071 ± 0.1240 0.5609 ± 0.0432
BIOTyang2023biot 0.7536 ± 0.0144 0.6004 ± 0.0195 0.6875 ± 0.0186 0.8761 ± 0.0284 0.3277 ± 0.0460 0.7068 ± 0.0457
LaBraMjianglarge 0.7721 ± 0.0093 0.5999 ± 0.0155 0.6909 ± 0.0125 0.8679 ± 0.0199 0.3287 ± 0.0402 0.7075 ± 0.0358
EEGPTwang2024eegpt 0.7162 ± 0.0171 0.5081 ± 0.0275 0.5597 ± 0.0171 0.8892 ± 0.0066 0.3073 ± 0.0641 0.5481 ± 0.0151
CBraModwang2024cbramod 0.7905 ± 0.0073 0.6267 ± 0.0099 0.7256 ± 0.0132 0.8892 ± 0.0154 0.3689 ± 0.0382 0.7398 ± 0.0284
@@
4.4 Ablation Study
We conduct ablation studies on three datasets using the same five seeds as in the main experiments to evaluate the key components in CodeBrain, as shown in Table 3. Additional results on mask ratio and architecture efficiency are in the supplementary material. Below is the detailed analysis:
(1) Tokenizer configuration: we compare the proposed TFDual-Tokenizer (Dual) with variants using only one domain codebook (Temporal or Frequency), or a shared codebook jointly reconstructing both domains (Mixed). Across all datasets, the Dual codebook consistently yields superior performance, suggesting that maintaining separate temporal and frequency codebooks captures complementary information more effectively than merging them.
(2) Sliding Window Attention (SWA): Including the SWA module consistently improves performance, confirming its regularization effect in capturing local dependencies. Notably, on the SEED-V dataset with 1s samples, SWA yields the largest gain, confirming its effectiveness for short-duration inputs.
(3) Contrastive learning (CL): We evaluate the impact of contrastive learning in TFDual-Tokenizer pretraining. It also leads to consistent gains, indicating a better capture of temporal patterns.
(4) Scaling Laws: Prior work wang2024cbramod ; jianglarge has explored scaling with 1 to 1000 hours of EEG data for pretraining and model sizes up to 5M parameters. We extend this by analyzing performance using 1k to 9k hours of pretraining data and larger models. As shown in Figure 3, Kappa, our primary evaluation metric, consistently improves with more data and parameters, suggesting that further scaling may lead to continued performance gains.

### v2

@@ 3.2 TFDual-Tokenizer Pretraining
where $\mathrm{sim}(e^{h}_{m1},e^{h}_{m2})$ is the cosine similarity, $\tau>0$ is the temperature parameter, and $\mathbb{1}_{[k\neq i]}$ is an indicator function used to exclude itself. We introduce a T-Decoder to reconstruct raw signals from the temporal code embedding $\tilde{e}_{i}$, which consists of a Transformer encoder followed by a linear projection. Let $y_{t}$ be the Transformer output from Eq. (2); the overall training objective is:
[…]
Finally, the training objective for the TFDual-Tokenizer can be defined as:
@@
$\mathcal{L}_{\text{tokenizer}}=\sum_{X_{m}\in\mathcal{X}}\mathcal{L}^{CL}_{m}+\sum_{X_{m}\in\mathcal{X}}\sum_{i=0}^{n}\ \mathcal{L}_{i}^{f}+\mathcal{L}^{t}_{i}$ $+\underbrace{\|\mathrm{sg}(\tilde{e}_{i})-vt_{pt_{i}}\|^{2}+\|\mathrm{sg}(\tilde{e}_{i})-vf_{pf_{i}}\|^{2}}_{\text{codebook loss}}$ (10)
$+\underbrace{\|\tilde{e}_{i}-\mathrm{sg}(vt_{pt_{i}})\|^{2}+\|\tilde{e}_{i}-\mathrm{sg}(vf_{pf_{i}})\|^{2}}_{\text{commitment loss}},$
@@
where $\mathrm{sg}(\cdot)$ denotes stop-gradient to avoid updating encoder parameters.
@@ 3.3 EEGSSM Pretraining
In this stage, we introduce a novel convolutional structured state space model framework, called EEGSSM, composed of multiple EEGSSM blocks. To adapt to unseen channels, we first learn dynamic positional embeddings that capture inter-channel relationships through a lightweight convolutional module. The resulting features are processed by EEGSSM blocks. Another 1D convolutional layer then maps the output back to the token space for reconstructing the indices of masked tokens produced by the TFDual-Tokenizer.
EEGSSM Block
Our EEGSSM block is composed of several blocks, which are integrated together through a residual connection mechanism. An EEGSSM block consists of a Layer Normalization, SGConv layer, SWA layer, and a gating component. Afterward, we feed the intermediate variables into the SGConv layer to obtain a global receptive field through convolution SSM.
SGConv Layer. SGConv is a structured SSM model (see Appendix B) using convolution architecture, and its convolution structure can be represented as a DFT formula:
@@
where $F_{N}$ denotes the DFT matrix of size $N$, and the convolution can be computed in $O(N\log N)$ via FFT. As a type of convolutional SSM, SGConv improves the convolution kernel $\overline{K}$ in Eq. (22) by introducing two features: sparse parameterization and kernel decay, making SGConv easier and more efficient to compute compared to the traditional S4 kernel. Let $L$ be the length of the input sequence. The convolution kernel $\overline{K}$ of SGConv is composed of several sub-kernels. Assuming the size of the first sub-kernel is $d$, with parameters $w_{i}\in\mathbb{R}^{d\times d}$, then the number of sub-kernels can be expressed as $N=\log_{2}(\frac{L}{d})+1$. The convolution kernel $\overline{K}$ in Eq. (22) can thus be initialized as:
@@
where $\alpha$ denotes the decay coefficient, usually chosen to be 0.5, inducing the decaying structure, and $\text{Upsample}(x)$ denotes upsample $x$ to length $l$. We also introduce the normalization constant $Z$ to ensure that the convolution operation does not change the scale of the input.
@@
Gate Mechanism. we use a gating mechanism to control the output of the block. We employ a gated unit similar to Wavenet (obeid2016temple, ), which can suppress useless or irrelevant features and help stabilize training in deep networks. We concatenate the global features output $y_{sg}$ by the SGConv layer with the local features output $y_{swa}$ by SWA and feed them into a gated unit:
@@
where $tanh(\cdot)$ and $\sigma(\cdot)$ are tanh function and sigmoid function, $\odot$ denotes an element-wise multiplication operator, $W_{f}$ and $W_{g}$ are learnable convolution filters. $y_{1}$ becomes the input to the next block, while $y_{2}$ will be aggregated to the output of SSM blocks through a skip connection.
Pre-Training Objective
To help the EEGSSM model learn general EEG representations, we use a Masked Autoencoder (MAE) for self-supervised pre-training. For a patched sample $X=\{x_{i}\mid i\in[1,2,\ldots,C]\}$, we randomly generate a mask $\mathcal{M}=\{m_{i}\mid i\in[1,2,\ldots,C]\}$ from a Bernoulli distribution of $r$ proportion, where $m_{i}\in\{0,1\}$. We reconstruct the token indices of masked EEG patches from the TFDual-Tokenizer by cross-entropy loss. Let $y_{i}$ denotes the output of the EEGSSM block, the probability that the EEG signal matches the corresponding token $v_{i}$ in the codebooks:
@@ 4.1 Datasets
Pre-Training. We pretrain CodeBrain on the TUH EEG Corpus (obeid2016temple, ), the largest publicly available EEG dataset to date. Data processing follows a standardized pipeline: recordings shorter than 5 minutes are excluded, and the first and last minute of each segment is removed. We retain 19 commonly used EEG channels (C3, C4, Cz, F3, F4, Fp1, Fp2, F7, F8, Fz, O1, O2, P3, P4, Pz, T3, T4, T5, T6), selected based on the international 10–20 system for electrode placement (hh1958ten, ). We apply band-pass filtering (0.3–75 Hz), and notch filtering at 60 Hz to remove noise. The data is resampled to 200 Hz and divided into 30-second non-overlapping segments. Segments with absolute amplitudes over $100\penalty 10000\ \mu\mathrm{V}$ are filtered out. To normalize the signals, each value is divided by 100. Each segment is split into 1-second windows, resulting in 570 EEG patches per sample. After preprocessing, 1,109,545 samples (about 9,246 hours) are retained for pretraining.
Downstream Tasks. We evaluate CodeBrain on 8 downstream tasks across 10 public EEG datasets, which span diverse applications and exhibit distribution shifts from the pretraining dataset, to assess generalizability. Detailed dataset configurations are in Table 1. We perform cross-subject or cross-session splits with strict separation between training, validation, and test sets. For FACED, we use 80 subjects for training, 20 for validation, and the remaining 23 for testing. In SEED-V, each session consists of 15 trials, which are evenly divided into training, validation, and test sets. ISRUC_S3 consists of 10 subjects, for which we apply an 8:1:1 cross-subject split. MentalArithmetic consists of 36 subjects, and we use a 7:1:1 cross-subject split. BCIC2020-T3 follows the official competition protocol. In CHB-MIT, we use recordings from 19 subjects for training, and from 2 subjects each for validation and testing. Additional dataset details are provided in the Appendix F.
@@
Downstream Tasks Datasets #Channels Length #Samples Class
Emotion Recognition FACED (chen2023large, ) 32 10s 10,332 9-class
SEED-V (liu2021comparing, ) 62 1s 117,744 5-class
Sleep Staging ISRUC_S1 (khalighi2016isruc, ) 6 30s 86,320 5-class
ISRUC_S3 (khalighi2016isruc, ) 6 30s 8,500 5-class
Imagined Speech Classification BCIC 2020-T3 (jeong20222020, ) 64 3s 6,000 5-class
Mental Stress Detection Mental Arithmetic (mumtaz2016eeg, ) 20 5s 1,707 2-class
Seizure Detection CHB-MIT (shoeb2009application, ) 16 10s 326,993 2-class
Motor Imagery Classification SHU-MI (goldberger2000physiobank, ) 32 4s 11,988 2-class
Event Type Classification TUEV (obeid2016temple, ) 16 5s 112,491 6-class
Abnormal Detection TUAB (obeid2016temple, ) 16 10s 409,455 2-class
@@
Cohen’s Kappa Weighted F1 Balanced Acc Cohen’s Kappa Weighted F1 Balanced Acc
BENDR 0.4716 ± 0.0095 0.5340 ± 0.0086 0.5320 ± 0.0083 0.0335 ± 0.0062 0.2026 ± 0.0330 0.2231 ± 0.0059
BIOT 0.4476 ± 0.0254 0.5136 ± 0.0112 0.5118 ± 0.0118 0.2261 ± 0.0262 0.3856 ± 0.0203 0.3837 ± 0.0187
LaBraM 0.4698 ± 0.0102 0.5288 ± 0.0188 0.5273 ± 0.0107 0.2386 ± 0.0209 0.3974 ± 0.0111 0.3976 ± 0.0138
EEGPT 0.4639 ± 0.0023 0.3924 ± 0.0017 0.4607 ± 0.0014 0.1323 ± 0.0062 0.3090 ± 0.0052 0.3061 ± 0.0044
CBraMod 0.5041 ± 0.0122 0.5618 ± 0.0093 0.5509 ± 0.0089 0.2569 ± 0.0143 0.4101 ± 0.0108 0.4091 ± 0.0097
CodeBrain 0.5406 ± 0.0084 0.5953 ±0.0113 0.5941 ± 0.0098 0.2735 ± 0.0032 0.4235 ± 0.0022 0.4137 ± 0.0023
@@
Cohen’s Kappa Weighted F1 Balanced Acc Cohen’s Kappa Weighted F1 Balanced Acc
BENDR 0.5995 ± 0.0151 0.6789 ± 0.0142 0.6352 ± 0.0095 0.0607 ± 0.0093 0.2379 ± 0.0165 0.2485 ± 0.0075
BIOT 0.7168 ± 0.0119 0.7834 ± 0.0096 0.7598 ± 0.0109 0.3650 ± 0.0176 0.4917 ± 0.0079 0.4920 ± 0.0086
LaBraM 0.7194 ± 0.0162 0.7843 ± 0.0189 0.7617 ± 0.0122 0.3800 ± 0.0242 0.5054 ± 0.0205 0.5060 ± 0.0155
EEGPT 0.6160± 0.0856 0.6375 ± 0.0632 0.6650 ± 0.0311 0.0567 ± 0.0164 0.2441 ± 0.0105 0.2453 ± 0.0131
CBraMod 0.7407 ± 0.0251 0.8056 ± 0.0219 0.7844 ± 0.0126 0.4216 ± 0.0163 0.5383 ± 0.0096 0.5373 ± 0.0108
@@
AUROC AUC-PR Balanced Acc AUROC AUC-PR Balanced Acc
BENDR 0.6248 ± 0.0765 0.3661 ± 0.0672 0.5681 ± 0.0448 0.8632 ± 0.0526 0.3071 ± 0.1240 0.5609 ± 0.0432
BIOT 0.7536 ± 0.0144 0.6004 ± 0.0195 0.6875 ± 0.0186 0.8761 ± 0.0284 0.3277 ± 0.0460 0.7068 ± 0.0457
LaBraM 0.7721 ± 0.0093 0.5999 ± 0.0155 0.6909 ± 0.0125 0.8679 ± 0.0199 0.3287 ± 0.0402 0.7075 ± 0.0358
EEGPT 0.7162 ± 0.0171 0.5081 ± 0.0275 0.5597 ± 0.0171 0.8892 ± 0.0066 0.3073 ± 0.0641 0.5481 ± 0.0151
CBraMod 0.7905 ± 0.0073 0.6267 ± 0.0099 0.7256 ± 0.0132 0.8892 ± 0.0154 0.3689 ± 0.0382 0.7398 ± 0.0284
@@
4.2 Experiment Settings
Experiment Setup. (1) Pretraining Setup. All experiments are conducted on NVIDIA 40GB A100 GPUs. The TFDual-Tokenizer is trained with temporal and frequency codebooks of 4096 codes (32 dimensions) for 20 epochs, and an 8-layer EEGSSM backbone with a masking ratio of 0.5 is trained for 10 epochs.
(2) Finetuning Strategy. We evaluate the quality of the pretrained representations under full finetuning. All downstream task datasets are resampled to 200 Hz to match the pretraining configuration. A three-layer MLP is applied to aggregate channel information, compress the $x$-second sequence, and map the representation to the target class, with activation and dropout between layers.
Baselines. We compare our model with five publicly available EFMs that have released pretrained weights, representing different pretraining strategies. BENDR (kostas2021bendr, ) adopts contrastive learning framework. BIOT (yang2023biot, ) uses patch-based continuous tokenization. LaBraM introduces discrete neural tokens through vector quantization (jianglarge, ). EEGPT (wang2024eegpt, ) and CBraMod (wang2024cbramod, ) rely on masked reconstruction of raw EEG signals.
More details about hyperparameter, baseline, and evaluation metric are provided in Appendix H, G.
4.3 Comparison with Baselines
We ensure consistent data splits across all baselines. Results are reported on six representative downstream datasets, with additional results provided in Appendix I. As shown in Table 2, CodeBrain achieves consistent performance gains compared to baselines. For multi-class classification, it achieves the largest gain of +0.0911 in Cohen’s Kappa (21.6%), +0.0718 in Weighted F1 score (13.3%), +0.0728 in Balanced Acc (13.5%) on BCIC 2020-T3 over the strongest baseline (wang2024cbramod, ). For binary classification, it achieves the largest gain of +0.0802 in AUROC (10.1%), +0.0910 in AUC-PR (14.5%) and +0.0258 in Balanced Acc (3.6%) on Mental Arithmetic over the strongest baseline. These results demonstrate the superior generalizability of CodeBrain.